3 ( t + 35 ) / 6 = 35
How do you work this out quickest possible way?!
Thankyou

Answer by nerdybill(7384)

3 ( t + 35 ) / 6 = 35
reduce the 3/6:
( t + 35 ) / 2 = 35
multiply both sides by 2:
t + 35 = 70
subtract 35 from both sides:
t = 35